# NonGEMM Bench: Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v3

@@ 4.3 Key Observations and Insights
  • Accelerating GEMM operators using GPUs significantly shifts the performance horizon, which makes the non-GEMM operators main bottleneck of the systems: On average, non-GEMM breakdown increases from 27% (CPU-only) to 55% (CPU + GPU).
  • We observe the dominance of non-GEMM operators after GPU acceleration across all the platforms (mobile, workstation, and datacenter) and task domains.
  • The choice of deployment software flow affects the most time-consuming non-GEMM operators: In ORT, memory operators dominate the execution time across all models (on average, 56%), while PyTorch doesn’t show clear patterns.
  • Dominant non-GEMM operator class vary depending on the task category: In IC models, normalization operators dominate (on average, 18.4% of the total execution time), while activation and element-wise arithmetic operators dominate in language models (17.75% and 17.6% for each, on average).
  • The profiling results indicate that the non-GEMM operators are now the prime optimization target after GEMM acceleration.
5 Related Works
Table 5: Comparing NonGEMM Bench with existing benchmarks: MLPerf [36], LongTail Bench [27], TorchBench [17], and NonGEMM Bench (This work).
Benchamrk MLPerf LongTailBench TorchBench NonGEMMBench (This work)
Real Usage Driven
Non-GEMM Focused
Real Dataset Driven
Plug Model & Profile
Based on the characteristics of non-GEMM operators, and on the results shown in  Section 4, we clarify different features of key related works comparing against this work in Table 5. TorchBench [17] presents a comprehensive benchmark suite for the Pytorch framework consisting of 84 models whose selection was driven by real usage like NonGEMM Bench. MLPerf [36] is an industry standard state-of-the-art inference benchmark, which focuses on the end-to-end performance. LongTail Bench [27] is a micro-benchmark of NonGEMM operators. However, LongTail bench is a microbenchmark that only focuses on custom implementations of 34 operations, which cannot accept new operators nor profile various implementation of the same operator. As summarized in Table 5, all work shed a light to individual important aspect, but only NonGEMM Bench provides all the desired features for a non-GEMM operator-focused benchmark.
6 Conclusion
Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law in favor of non-GEMM operators. To guide future optimization, we need to understand the performance of non-GEMM operators in recent popular models. In this work, we analyze the performance non-GEMM operators in recent popular machine learning models. We construct NonGEMM Bench to study of non-GEMM operators in accelerated popular models reflecting realistic use scenarios and evaluate it on different hardware platforms. We show that the dominance of non-GEMM operators is occurring in all platforms with GPUs, and the deployment software’s impact is large. Our findings strongly motivate ML system researchers to consider non-GEMM operators as a major consideration.

### v4

@@ IV-D Key Observations and Insights
  • After GEMM acceleration, non-GEMM becomes a major optimization target regardless of the hardware platforms.
  • Specialized optimization for one non-GEMM is not effective due to the diversity in dominant non-GEMM.
  • Operator fusion cannnot fully address non-GEMM performance challenge: Although it can significantly improve non-GEMM performance, but its effectiveness heavily depends on the model.
  • Operator support in deployment flows significantly affects the non-GEMM performance horizon.
  • Quantization significantly aggravates the non-GEMM performance challenge due to the imbalanced speedup across GEMM and non-GEMM and quantization/dequantization costs around non-GEMM operators
V Related Works
Previous works characterized ML workloads and non-GEMM operators. However, the main limitations are listed below: State-of-the-art inference benchmarks  [20, 52] do not thoroughly analyze the performance of Non-GEMM operators. Other works characterizing non-GEMM operators focus on a specific application domain [48, 47, 65, 34] for their non-GEMM study. In [48],  [27], and  [47] the authors analyze the performance of non-GEMM operators, however they focus on a specific class of transformer language models. Similarly, [34] only studies a limited number of operators from computer vision models in a microbenchmarking flow. Previous works that also studied non-GEMM operator performance bottlenecks [12, 56, 60, 51], mostly focused on softmax operators in transformers for language models. Our work extends on the pioneering efforts of these works by studying non-GEMM operators from 15 models covering 4 various task domains.
VI Conclusion
Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law to have more presence of non-GEMM operators in the performance horizon. In this paper, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms. The results confirm the increasing importance of non-GEMM performance and show model optimizations we apply (e.g., quantization) can significantly aggravate the non-GEMM performance challenge. The dominant non-GEMM operators are diverse across models, which indicates that a specialized optimization targeting a specific operator cannot solve the non-GEMM performance challenge. We also show that non-GEMM-oriented optimization such as operator fusion cannot fully address the non-GEMM performance challenge.
Our performance anlysis results imply that now we need to consider non-GEMM operators as a major optimization target and develop new hardware and software techniques to optimize non-GEMM performance. The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to such follow-up research.